OpenStudy (rmrjr22):

integral 1/(x^2-x+1) dx Steps?

OpenStudy (amistre64):

complete the square maybe

myininaya (myininaya):

Trig sub for the win Need to write bottom as the difference of two squares

myininaya (myininaya):

and let me say "squares"

OpenStudy (amistre64):

the bottom is NOT a diff of 2 squares tho is it?

OpenStudy (experimentx):

x^2 - 2*(1/2) x + (1/4) +(1-1/4) => (x-1/2)^2 + (sqrt(3)/2)^2

OpenStudy (amistre64):

\[\frac{1}{(x-\frac{1}{2})^2+\frac{3}{4}}\]

myininaya (myininaya):

lol... \[x^2-x+1=x^2-x+(\frac{1}{2})^2+1-(\frac{1}{2})^2=(x-\frac{1}{2})^2+1-\frac{1}{4}=(x-\frac{1}{2})^2+\frac{3}{4}\] \[=(x-\frac{1}{2})^2+(\frac{\sqrt{3}}{2})^2\] lol yes Or i mean the sum of two squares pardon my earlier french

myininaya (myininaya):

or one of the two depending on what you gots

OpenStudy (rmrjr22):

how does the (1/2)^2 come into the equation?

OpenStudy (amistre64):

\[\frac{1}{(x-\frac{1}{2})^2+\frac{3}{4}}\frac{4/3}{4/3}=\frac{4/3}{\frac{4}{3}(x-\frac{1}{2})^2+1} \] i dont spose theres a way to get this into the tan-1 stuff ...

OpenStudy (amistre64):

partial decomp in the complexes i spose

OpenStudy (rmrjr22):

Thanks!

OpenStudy (amistre64):

yw, hope it helps